# Enhancing Quantum Key Distribution with Entanglement Distillation and Classical Advantage Distillation

### v1

@@ Abstract
Realizing secure communication between distant parties is one of quantum technology’s main goals. Although quantum key distribution promises information-theoretic security for sharing a secret key, the key rate heavily depends on the level of noise in the quantum channel. To overcome the noise, both quantum and classical techniques exist, i.e., entanglement distillation and classical advantage distillation. So far, these techniques have only been used separately from each other. Herein, we present a two-stage distillation scheme concatenating entanglement distillation with classical advantage distillation. While for the advantage distillation, we use a fixed protocol, i.e., the repetition code, in the case of entanglement distillation, we employ an enumeration algorithm to find the optimal protocol. We test our scheme for different noisy entangled states and demonstrate its quantitative advantage: our two-stage distillation scheme achieves finite key rates even in the high-noise regime where entanglement distillation or advantage distillation alone cannot afford key sharing. Since the advantage distillation part does not introduce further requirements on quantum resources, the proposed scheme is well-suited for near-term quantum key distribution tasks.
@@ I Introduction
The security of QKD protocols typically can be related to the presence of entanglement [2]. The intuition is that if both parties share (pure) maximally entangled states, then they can obtain correlated outcomes that are independent of any third party, which renders any eavesdropping attack impossible. However, in the presence of realistic devices quantum information processing is necessarily noisy. This noise degrades the rate at which a quantum key distribution protocol produces secure keys [3].
@@
In order to overcome noise, there are several possible strategies. If the key is produced by measuring an entangled state, entanglement distillation (ED) [4, 5] can transform probabilistically a number of noisy entangled states into a smaller number of more entangled states. Entanglement distillation has already been demonstrated in experiments, e.g. [6, 7, 8, 9]. However, state-of-the-art quantum communication hardware can only store a handful of states, limiting the amount of entanglement distillation possible.
@@
Alternatively, it is possible to act directly on the classical outcomes obtained by measuring the quantum resources. Advantage distillation (AD) [10, 11] is a purely classical procedure that transforms probabilistically a bit string on which an eavesdropper has a large amount of information into a smaller bit string on which the eavesdropper has less information. While there is no limitation on the amount of advantage distillation that the legitimate key distribution parties can apply, it is a less efficient process than entanglement distillation with fundamental differences in the amount of noise that it can overcome [12].
@@
Herein, we propose a state-aware two-stage key distillation scheme by concatenating entanglement distillation with advantage distillation. An obvious advantage of this scheme is that it does not require additional quantum resources than those required to perform the entanglement distillation in the first stage alone. Our scheme builds on a recent enumeration of entanglement distillation algorithms [13, 14] which allows us to optimize the two-stage scheme for a given starting state. In particular, by recognizing the equivalence of an advantage distillation protocol and a CNOT-based entanglement distillation protocol, we generalize the enumeration scheme to concatenated entanglement distillation with advantage distillation (ED+AD) protocols. We quantitatively demonstrate that this two-stage distillation scheme can achieve non-zero key rates at a heavy noise regime, whereas using either stage alone could not achieve positive key rates.
@@
The paper is structured as follows: in Sec. II.1, we review the basics of entanglement-based quantum key distribution. In Sec. II.2, we introduce the mathematical structure of bi-local Clifford distillation protocols, which form an important subset of entanglement distillation protocols. Some mathematical details will be provided in the appendix. In Sec. II.3, we review the idea of classical advantage distillation and discuss its connection to bi-local Clifford distillation protocols. In Sec. III, we describe the concatenated ED+AD protocols and the method we used to enumerate all possible such protocols for the optimization of QKD performance. In the results section, we present and discuss the quantitative performance of the ED+AD protocols for two particular noisy entangled states.
@@ II Background
Here we review the technical background on entanglement-based key distribution, and two main techniques employed in this paper, including bi-local Clifford entanglement distillation and classical advantage distillation.
@@ II.1 Entanglement-based quantum key distribution
Quantum key distribution protocols can be categorized into the prepare-and-measure schemes (e.g., the original form of the BB84 protocol [1]) or the entanglement-based schemes (e.g., the E91 protocol [2]). However, it has been shown that these two categories are mathematically equivalent [15]. For example, the prepare-and-measure BB84 protocol can be performed in the following way based on pre-shared entanglement.
@@
6:Information reconciliation and privacy amplification [16] are performed to obtain secure keys.
@@ II.2 Bi-local Clifford entanglement distillation protocols
Entanglement distillation refers to turning a number $m$ of noisy entangled pairs into a smaller number $k$ of more entangled pairs by means of local quantum operations and classical communication [4]. Due to their mathematical structure, the bi-local Clifford entanglement distillation protocols have been extensively studied [13, 14, 18]. Here we restrict our attention to bi-local Clifford protocols taking $m$ pairs to 1 ($m$-1 protocol), see Fig. 1 for an illustration of a $4$-$1$ bi-local Clifford entanglement distillation protocol. More precisely, an $m$-$1$ bi-local Clifford entanglement distillation protocol is defined as follows:
@@
1:Alice and Bob first share $m$ entangled pairs in the form of $|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$.
@@
Figure 1: Illustration of a bi-local Clifford entanglement distillation protocol. Alice and Bob first share some noisy entangled pairs. They then perform local Clifford operations ($C^{T}$ and $C^{\dagger}$) on their share of qubits. They measure all pairs except the first pair. If the measurement outcomes of the corresponding entangled pairs coincide, they keep the first pair (unmeasured one). Otherwise, they discard it.
There are many well-known protocols that belong to the bi-local Clifford class. For instance, DEJMPS protocol (also known as Oxford protocol) [5] is an example of a $2$-$1$ bi-local Clifford protocol. In DEJMPS the gate is given by $C^{T}=\left(\text{CNOT}_{12}\right)\left(R_{x,1}(\frac{\pi}{2})\otimes R_{x,2}(\frac{\pi}{2})\right)$; where the $\text{CNOT}_{12}$ is performed on local qubit pairs and $R_{x,i}$ is an x-axis rotation on the $i$-th qubit.
Bi-local Clifford protocols are a very general class of protocols. Optimizing over this class can be done by enumerating all elements of the Clifford group. However, the number of Clifford gates grows super exponentially with the number of qubits [13]. Fortunately, it is possible to simplify the enumeration by building on the mathematical correspondence between Clifford operations acting on the $n$-qubit Pauli group (ignoring the phase information) and the symplectic group $\textrm{Sp}(2n,\mathbb{Z}_{2})$ with its natural group action (readers are referred to appendix A for the details). In particular, Clifford gates $C_{1}$ and $C_{2}$, when used to perform the entanglement distillation could lead to the same distillation statistics (i.e., output fidelity and success probability) [13, 19]. This observation enables enumerating all different bi-local Clifford protocols by enumerating an element in each distillation equivalence class. This reduction allows an enumeration algorithm to find the optimal protocol for a small number of shared noisy entangled pairs.
@@ II.3 Advantage distillation and its relation to entanglement distillation
Advantage distillation (AD) is a well-known technique in secret key distillation scenarios. It allows Alice and Bob to enhance the correlation between their shared random variable and reduce its correlation to a third party [10]. It is designed so that Alice and Bob can gain an “advantage” over the eavesdropper, Eve, by recognizing and keeping the part of the shared random variables that has a higher correlation between Alice and Bob than Eve. Advantage distillation typically uses parity calculation and two-way communication to decide if a certain part of the key is kept or not [12, 20].
Let $A^{N}$ and $B^{N}$ be $N$-bit correlated bit strings. An example of an AD protocol works by Alice generating a random bit $s$, adding it bitwise to a block of $m$ bits ($a^{m}$) to obtain $C$ (which contains parity information of the block and the random bit). She then publicly sends $C$ to Bob via an authenticated channel. Bob calculates the bitwise addition of $C$ and his corresponding block ($b^{m}$). If the results are all-00 or all-$1$, Bob sends an accept signal back to Alice, and a correlated bit is obtained. The process is repeated several rounds until a key string with a suitable size is obtained.
@@
III ED+AD key distribution protocols
For entanglement-based quantum key distribution, both entanglement distillation and advantage distillation could increase the correlation of the shared key string. However, the former requires coherent quantum operations on the entangled states, while the latter can be performed after the quantum measurement and with the help of classical communication. Intuitively, entanglement distillation has the potential to increase the output fidelity with respect to a particular Bell state ($|\Phi^{+}\rangle$, in this paper) by correcting both bit-flip and phase-flip errors. On the other hand, the advantage distillation in the computational basis aims to correct the bit-flip error only.
Herein, we describe the structure of an $m$-$n$-$1$ ED+AD concatenating protocol for entanglement-based quantum key distribution in Protocol 3.
Note that the entanglement distillation stages as shown in Fig. 2 do not have to be successful at the same time. Since the advantage distillation is performed offline based on the measurement outcome, the individual entanglement distillation steps can be executed in a repeat-until-success manner.
Protocol 3 $m$-$n$-$1$ ED+AD key distribution protocol.
@@ III ED+AD key distribution protocols
Figure 2: Illustration of the ED+AD protocol. The noisy entangled pairs shared by Alice and Bob first undergo entanglement distillation. The successfully distilled pairs are measured. The measurement outcomes (dashed lines) become the input to the next advantage distillation protocol. The illustration corresponds to a 4-2-1 protocol, e.g. a concatenation of a 4-1 entanglement distillation protocol with a 2-1 advantage distillation protocol.
@@
Table 1: Distillation protocols and their constituents. In this paper, we studied 9 different distillation protocols. Their constituent protocols are summarized in this table.
One key point for concatenating two distillation protocols is that the advantage distillation is performed after the quantum measurement. This means that the ED+AD protocol does not require additional quantum resources than those required for performing entanglement distillation alone. However, we note the following equivalence between advantage distillation protocols and a CNOT-only Clifford distillation protocol. As shown in Fig. 3, since the CNOTs performed in the computational basis commute with the measurement, we can postpone measurement to the end and use the quantum state before the measurement to calculate relevant figures of merits for performing QKD protocols. This equivalence also allows us to view an ED+AD protocol as an $mn$-$1$ distillation without having to use an $mn$-qubit quantum computer.
@@
5:     for Single-qubit Pauli operator $P$ do Calculate the distilled state $\rho_{\text{distilled}}$ by concatenating distillation protocols $C_{m}$, Pauli operator $P$, and $A_{n}$.
@@ IV Results
In this section, we analyze the quantitative performance of the ED+AD $m$-$n$-1 protocols for two standard noise models. We enumerated all possible ED+AD protocols to find the optimal protocol for the two specific noisy states that result from applying the corresponding noise model. In appendix B, we also show that the optimal protocol found by the enumeration performs better than naively using the DEJMPS protocol in the entanglement distillation stage.
@@ IV.1 Depolarizing noise
We first analyze the performance of the ED+AD protocols in the case of the Werner input state (Eqn. IV.1[22], which results from applying depolarizing noise on the Bell state $|\Phi^{+}\rangle$. The Werner state is a highly symmetric quantum state that is parametrized by a single parameter (i.e., the input fidelity $F$). The depolarizing noise can be seen as a worst-case noise model from a standard twirling argument [4].
@@ IV.1.1 Achievable fidelity
The $m$-$n$-1 ED+AD protocols consistently perform better than their $m$-1 ED-only counterparts. It is interesting that in the heavy noisy regime, the protocol $3$-$2$-$1$ achieves a higher fidelity than the $2$-$3$-$1$ protocol. This implies that using the more flexible $3$-$1$ entanglement distillation protocols in the first ED stage is more advantageous in terms of producing more entangled distilled states. In contrast, the $2$-$4$-$1$ protocol achieves worse fidelity than the $2$-$3$-$1$ and $2$-$2$-$1$ protocols. This rather counter-intuitive behavior has its origin in the fact that, unlike entanglement distillation, the advantage distillation can only correct bit-flip errors in the computational basis. Therefore, the phase error component $|\Phi^{-}\rangle$ can be amplified during the distillation. This phenomenon is discussed in more detail in appendix C, where we provide analytical expressions of performing AD-only. Note that for a fidelity of $\frac{1}{2}$ or less the Werner state is separable, and no entanglement can be distilled.

### v2

@@ Abstract
Realizing secure communication between distant parties is one of quantum technology’s main goals. Although quantum key distribution promises information-theoretic security for sharing a secret key, the key rate heavily depends on the level of noise in the quantum channel. To overcome the noise, both quantum and classical techniques exist, i.e., entanglement distillation and classical advantage distillation. So far, these techniques have only been used separately from each other. Herein, we present a two-stage distillation scheme concatenating entanglement distillation with classical advantage distillation. For advantage distillation, we utilize a fixed protocol, specifically, the repetition code; in the case of entanglement distillation, we employ an enumeration algorithm to find the optimal protocol. We test our scheme for different noisy entangled states and demonstrate its quantitative advantage: our two-stage distillation scheme achieves finite key rates even in the high-noise regime where entanglement distillation or advantage distillation alone cannot afford key sharing. We also calculate the security bounds for relevant QKD protocols with our key distillation scheme and show that they exceed the previous security bounds with only advantage distillation. Since the advantage distillation part does not introduce further requirements on quantum resources, the proposed scheme is well-suited for near-term quantum key distribution tasks.
@@ I Introduction
Typically, the security of QKD protocols can be related to the presence of entanglement [2]. The intuition is that if both parties share (pure) maximally entangled states, they can then obtain correlated outcomes that are independent of any third party, which renders any eavesdropping attack impossible. However, in the presence of realistic devices, quantum information processing is necessarily noisy. This noise degrades the rate at which a quantum key distribution protocol produces secure keys [3].
@@
There are several possible strategies to overcome noise. If the key is produced by measuring an entangled state, entanglement distillation (ED) [4, 5] can probabilistically transform a number of noisy entangled states into a smaller number of more entangled states. Entanglement distillation has already been demonstrated in experiments, e.g., [6, 7, 8, 9, 10]. However, state-of-the-art quantum communication hardware can only store a handful of states, limiting the amount of entanglement distillation possible.
@@
Alternatively, it is possible to act directly on the classical outcomes obtained by measuring the quantum resources. Advantage distillation (AD) [11, 12] is a purely classical procedure that transforms a bit string probabilistically on which an eavesdropper has a large amount of information into a smaller bit string on which the eavesdropper has less information. While there is no limitation on the amount of advantage distillation that the legitimate key distribution parties can apply, it is a less efficient process than entanglement distillation with fundamental differences in the amount of noise that it can overcome [13, 14].
@@
Herein, we propose a state-aware two-stage key distillation scheme by concatenating entanglement distillation with advantage distillation. An obvious advantage of this scheme is that it does not require additional quantum resources than those required to perform the entanglement distillation in the first stage alone. Our scheme builds on a recent enumeration of entanglement distillation algorithms [15, 16], which allows us to optimize the two-stage scheme for a given starting state. In particular, by recognizing the equivalence of an advantage distillation protocol and a CNOT-based entanglement distillation protocol, we generalize the enumeration scheme to concatenated entanglement distillation with advantage distillation (ED+AD) protocols. We quantitatively demonstrate that this two-stage distillation scheme can achieve non-zero key rates at a heavy noise regime, whereas using either stage alone could not achieve positive key rates.
@@
The paper is structured as follows: in Sec. II.1, we review the basics of entanglement-based quantum key distribution. In Sec. II.2, we introduce the mathematical structure of bi-local Clifford distillation protocols, which form an important subset of entanglement distillation protocols; further mathematical details will be provided in the appendix. In Sec. II.3, we review the idea of classical advantage distillation and discuss its connection to bi-local Clifford distillation protocols. In Sec. III, we describe the concatenated ED+AD protocols and the method we used to enumerate all possible such protocols for the optimization of QKD performance. In the results section, we present and discuss the quantitative performance of the ED+AD protocols for two particular noisy entangled states. Importantly, we derive the critical quantum bit error rates (QBER) for two standard QKD protocols with the ED+AD key distillation scheme. We demonstrate the ED+AD strategy strictly tolerates more noise than using advantage distillation alone.
@@ II Background
Here, we review the technical background of entanglement-based key distribution and two main techniques employed in this paper, namely bi-local Clifford entanglement distillation and classical advantage distillation.
@@ II.1 Entanglement-based quantum key distribution
Quantum key distribution protocols can be categorized into the prepare-and-measure schemes (e.g., the original form of the BB84 protocol [1] and the six-state protocol [17]) or the entanglement-based schemes (e.g., the E91 protocol [2]). However, it has been shown that these two categories are mathematically equivalent [18]. For example, the prepare-and-measure BB84 protocol can be performed in the following way based on pre-shared entanglement.
@@
6:Information reconciliation and privacy amplification are performed to obtain secure keys. [19]
@@ II.2 Bi-local Clifford entanglement distillation protocols
Entanglement distillation refers to turning a number $m$ of noisy entangled pairs into a smaller number $k$ of more entangled pairs by means of local quantum operations and classical communication [4]. Due to their mathematical structure, the bi-local Clifford entanglement distillation protocols have been extensively studied [15, 16, 21]. Here, we restrict our attention to bi-local Clifford protocols taking $m$ pairs to 1 pair ($m$-1 protocol); see Fig. 1 for an illustration of a $4$-$1$ bi-local Clifford entanglement distillation protocol. More precisely, an $m$-$1$ bi-local Clifford entanglement distillation protocol is defined as follows:
@@
1:Alice and Bob first share $m$ entangled pairs in the form of $|\Phi^{+}\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|11\rangle\right)$ through noisy quantum channels.
@@
Figure 1: Illustration of a bi-local Clifford entanglement distillation protocol. Alice and Bob first share some noisy entangled pairs. They then perform local Clifford operations ($C^{T}$ and $C^{\dagger}$) on their share of qubits. They measure all pairs except the first pair. If the measurement outcomes of the corresponding entangled pairs coincide, they keep the first unmeasured pair. Otherwise, they discard it.
There are many well-known protocols that belong to the bi-local Clifford class. For instance, the DEJMPS protocol (also known as Oxford protocol) [5] is an example of a $2$-$1$ bi-local Clifford protocol. In DEJMPS, the Clifford operation is given by $C^{T}=\left(\text{CNOT}_{12}\right)\left(R_{x,1}(\frac{\pi}{2})\otimes R_{x,2}(\frac{\pi}{2})\right)$; where the $\text{CNOT}_{12}$ is performed on local qubit pairs and $R_{x,i}$ is an x-axis rotation on the $i$-th qubit.
Bi-local Clifford protocols are a very general class of protocols. Optimizing over this class can be done by enumerating all elements of the Clifford group. However, the number of Clifford gates grows super exponentially with the number of qubits [15]. Fortunately, it is possible to simplify the enumeration by building on the mathematical correspondence between Clifford operations acting on the $n$-qubit Pauli group (ignoring the phase information) and the symplectic group $\textrm{Sp}(2n,\mathbb{Z}_{2})$ with its natural group action (readers are referred to appendix A for the details). In particular, Clifford gates $C_{1}$ and $C_{2}$, when used to perform the entanglement distillation, could lead to the same distillation statistics (i.e., output fidelity and success probability) [15, 22]. This observation enables enumerating all different bi-local Clifford protocols by enumerating an element in each distillation equivalence class. This reduction allows an enumeration algorithm to find the optimal protocol for a small number of shared noisy entangled pairs.
@@ II.3 Advantage distillation and its relation to entanglement distillation
Advantage distillation (AD) is a well-known technique in secret key distillation scenarios. It allows Alice and Bob to enhance the correlation between their shared random variable and reduce its correlation to a third party [11]. It is designed so that Alice and Bob can gain an “advantage” over the eavesdropper—Eve—by recognizing and keeping the part of the shared random variables that has a higher correlation between Alice and Bob than Eve. Advantage distillation typically uses parity calculation and two-way communication to decide if a certain part of the key is kept or not [14, 23].
We now provide an example of an AD protocol. Let $A^{N}$ and $B^{N}$ be $N$-bit correlated bit strings belonging to Alice and Bob, respectively. Alice generates a random bit $s$ and adds it bitwise to a block of $m\leq N$ bits ($a^{m}$) to obtain $C$ (which contains parity information of the block and the random bit). She then publicly sends $C$ to Bob via an authenticated channel. Bob calculates the bitwise addition of $C$ and his corresponding block ($b^{m}$). If the results are all-00 or all-$1$, Bob sends an accept signal back to Alice, and a correlated bit is obtained. The process is repeated several rounds until a key string with a suitable size is obtained.
@@
III ED+AD key distillation protocols
For entanglement-based quantum key distribution, both entanglement distillation and advantage distillation increase the correlation of the shared key string. However, the former requires coherent quantum operations on the entangled states, while the latter can be performed after the quantum measurement and with the help of classical communication. Intuitively, entanglement distillation has the potential to increase the output fidelity with respect to a particular Bell state ($|\Phi^{+}\rangle$, in this paper) by correcting both bit-flip and phase-flip errors. On the other hand, the advantage distillation in the computational basis aims to correct the bit-flip error only.
Herein, we describe the structure of an $m$-$n$-$1$ ED+AD concatenated protocol for entanglement-based quantum key distribution in Protocol 3.
Note that the entanglement distillation stages, as shown in Fig. 2, do not have to be successful at the same time. Since the advantage distillation is performed offline based on the measurement outcome, the individual entanglement distillation steps can be executed in a repeat-until-success manner.
Protocol 3 $m$-$n$-$1$ ED+AD key distillation protocol.
@@ III ED+AD key distillation protocols
Figure 2: Illustration of the ED+AD protocol. The noisy entangled pairs shared by Alice and Bob first undergo entanglement distillation. The successfully distilled pairs are measured. The measurement outcomes (dashed lines) become the input to the advantage distillation protocol. The illustration corresponds to a 4-2-1 protocol, e.g., a concatenation of a 4-1 entanglement distillation protocol with a 2-1 advantage distillation protocol.
@@
Table 1: Distillation protocols and their constituents. In this paper, we studied nine different distillation protocols. Their constituent protocols are summarized in this table.
One key point for concatenating two distillation protocols is that the advantage distillation is performed after the quantum measurement. This means that the ED+AD protocol does not require additional quantum resources than those required for performing entanglement distillation alone. However, we note the following equivalence between advantage distillation protocols and a CNOT-only Clifford distillation protocol. As shown in Fig. 3, since the CNOTs performed in the computational basis commute with measurements, we can postpone the measurements to the end and use the quantum state before the measurement to calculate relevant figures of merit for QKD protocols. This equivalence also allows us to view an ED+AD protocol as an $mn$-$1$ distillation without having to use an $mn$-qubit quantum computer.
@@
5:     for Single-qubit Pauli operator $P$ do Calculate the distilled state $\rho_{\text{distilled}}$ by concatenating $C_{m}$, $P$, and $A_{n}$.
@@ IV Results
In this section, we analyze the quantitative performance of the ED+AD $m$-$n$-1 protocols for two standard noise models. We enumerated all possible ED protocols in combination with the repetition-code AD protocol to find the optimal ED+AD protocol for the two specific noisy states that result from applying the corresponding noise model. In appendix C, we also show that the optimal protocol found by the enumeration performs better than naively using the DEJMPS protocol in the entanglement distillation stage. Finally, we discuss the security bounds for the ED+AD key distillation scheme and show that they outperform previous security bounds where only advantage distillation is considered.
@@ IV.1 Depolarizing noise
We first analyze the performance of the ED+AD protocols in the case of the Werner input state (Eqn. IV.1[25], which results from applying depolarizing noise on the Bell state $|\Phi^{+}\rangle$. The definition of the Bell states is given in Eqn. A. The Werner state is a highly symmetric quantum state that is parametrized by a single parameter (i.e., the input fidelity $F$). The depolarizing noise can be seen as a worst-case noise model from a standard twirling argument [4].
@@ IV.1.1 Achievable fidelity
The $m$-$n$-1 ED+AD protocols consistently perform better than their $m$-1 ED-only counterparts. It is interesting that the protocol $3$-$2$-$1$ achieves a higher fidelity than the $2$-$3$-$1$ protocol. This implies that using the more flexible $3$-$1$ entanglement distillation protocols in the first ED stage is more advantageous in terms of producing more entangled distilled states. Unexpectedly, the $2$-$4$-$1$ protocol achieves worse fidelity than the $2$-$3$-$1$ and $2$-$2$-$1$ protocols. This rather counter-intuitive behavior has its origin in the fact that, unlike entanglement distillation, advantage distillation only corrects for bit-flip errors in the computational basis. Therefore, the phase error component $|\Phi^{-}\rangle$ can be amplified during the distillation. This phenomenon is discussed in more detail in appendix D, where we provide analytical expressions of performing AD-only. Note that for a fidelity of $\frac{1}{2}$ or less, the Werner state is separable, and no entanglement can be distilled.